OpenStudy (anonymous):

A quiz consists of 10 questions. Each question is marked either “right” or “wrong.” For each student in the class, the instructor keeps a record of the number of right answers the student got. In the class, the average number of right answers is 6.2 and the SD of the number of right answers is 1.4 find average of wrong answers

Directrix (directrix):

@salman_09 Would you read the following and determine what you think about it? Thanks. Insight from @Phone_A_Friend_SS_from_California "I think the answer is 1.4. If x is the number right score than 10-x would be the number wrong score. If M is the st. dev. of right scores then 10-M is the mean of the wrong scores. (assuming here that any non-right response, including leaving answer space blank, is a wrong score. The sd. dev. formula for number right involves summing (x-M)^2. The sd. dev. formula for number wrong scores involves summing [(10-x) - (10-M)]^2 = (-x+M)^2 = (x-M)^2."

Directrix (directrix):

http://openstudy.com/study#/updates/514015c6e4b0f08cbdc8ce5a Courtesy of @kropot72